OpenStudy (anonymous):

Find the x-intercept and the y-intercept for the graph of the equation. x=2y=10

OpenStudy (gorv):

is it x+2y=10??

OpenStudy (gorv):

for x intercept put y=0

OpenStudy (gorv):

@teabby

OpenStudy (gorv):

x+0=10 x=10 is x intercept

OpenStudy (gorv):

fro y intercept put x=0 0+2y=10

OpenStudy (gorv):

divide both side by 2

OpenStudy (gorv):

y=5 is y intercept

OpenStudy (anonymous):

thanks i just checked it thats right

OpenStudy (gorv):

yeah u got the ans now u will rply man..i was helpin u so it was ur duty to rply when i asked u

OpenStudy (anonymous):

i dont have to respond when you tell me to tf & i was doing the same thing you were doing working the problem out. i will reply when i reply.